# Spiral of Silence: How is Large Language Model Killing Information Retrieval?—A Case Study on Open Domain Question Answering

### v2

@@ Abstract
The practice of Retrieval-Augmented Generation (RAG), which integrates Large Language Models (LLMs) with retrieval systems, has become increasingly prevalent. However, the repercussions of LLM-derived content infiltrating the web and influencing the retrieval-generation feedback loop are largely uncharted territories. In this study, we construct and iteratively run a simulation pipeline to deeply investigate the short-term and long-term effects of LLM text on RAG systems. Taking the trending Open Domain Question Answering (ODQA) task as a point of entry, our findings reveal a potential digital "Spiral of Silence" effect, with LLM-generated text consistently outperforming human-authored content in search rankings, thereby diminishing the presence and impact of human contributions online. This trend risks creating an imbalanced information ecosystem, where the unchecked proliferation of erroneous LLM-generated content may result in the marginalization of accurate information. We urge the academic community to take heed of this potential issue, ensuring a diverse and authentic digital information landscape. [^1] We release the resources at https://github.com/VerdureChen/SOS-Retrieval-Loop
@@ 1 Introduction
Figure 1: The evolution of RAG systems after introducing LLM-generated texts, where the "Spiral of Silence" effect gradually emerges.
@@
Further examination reveals a bias in search systems towards LLM-generated texts, which consistently rank higher than human-written content. As LLM-generated texts increasingly dominate the search results, the visibility and influence of human-authored web content diminish, fostering a digital "Spiral of Silence" effect. This effect aptly explains what we observed in our simulations and reveals the potential negative impact of LLM-generated texts on the information ecosystem: while LLM-generated texts sometimes provide a more effective IR experience in the short term, in the long term they may lead to the invisibility of human-authored content, the homogenization of search results, and the inaccessibility of certain accurate information, thereby adversely affecting public knowledge acquisition and decision-making.
The contributions of this paper are threefold: 1) An investigation of the short-term and long-term impacts of LLM-generated text on RAG systems. 2) By performing an iterative ODQA pipeline, we study the potential emergence of a "Spiral of Silence" phenomenon within RAG systems. 3) We analyze the implications of this phenomenon, offering a new perspective on the dynamic interplay between LLM-generated content and RAG systems.
@@ 2 Related Works
Spiral of Silence. The "Spiral of Silence" theory (Noelle-Neumann, 1974), is a seminal theory within the field of communications that describes how people may suppress their views to avoid isolation, thus often reinforcing dominant public opinions (Scheufle and Moy, 2000; Liu et al., 2019; Lin et al., 2022). We shift focus to a novel "passive human silence" influenced by LLMs, where rapid AI content production and biased search algorithms potentially marginalize human contributions in public discourse. This theory stands apart from concepts such as "echo chambers", "filter bubbles", and "degenerate feedback loops" prevalent in recommendation systems  (Alatawi et al., 2021; Chitra and Musco, 2020; Jiang et al., 2019). While these terms describe the narrowing of informational scope as users engage with algorithmic systems, the "Spiral of Silence" theory proposes a scenario where human users become thoroughly passive—they fall silent in public discourse due to the influence of LLMs and the IR systems, which is not merely a result of selective exposure or algorithmic recommendations. Our study explores how LLM-generated text might induce a "Spiral of Silence" in RAG systems over time. For further discussion regarding the rationale for applying this theory to our study, please see Appendix A.1.
@@ 3.1 Preliminaries
Generation Stage, executed by the generation function $G:(P\times Q\times D^{\prime}\times K)\rightarrow S$. Its task is to utilize the prompt $p\in P$, the query $q$, the related document subset $D^{\prime}$, and the knowledge of LLMs $K$ to construct the answer $s$.
@@ 3.2 Simulation Process
Our simulation process starts with a pure human-authored text dataset and gradually introduces the LLM-generated text, observing how this change over time affects the RAG system. Adhering to the specifications outlined in Section 3.1, the RAG architecture is instantiated and expanded with additional details. In the retrieval stage, we apply sparse and dense retrieval strategies to obtain a candidate document set that is relevant to the query. Additionally, we also have the option to perform a re-ranking of the candidate documents to further optimize the process. In the generation stage, we use the LLMs which are widely used in research and practical applications to generate responses. To accurately simulate the evolution process of the RAG system, we specifically use an iteratively updated indexing structure that supports incorporating the newly generated LLM text into the index in each iteration, keeping the dataset updated for subsequent retrieval and evaluation.
@@
The pseudo-code for this process is presented in Appendix A.2. Through the simulation process, we can observe how LLM-generated text influences the entire pipeline of RAG systems, and how this impact evolves over time and with data accumulation. Due to the relatively infrequent update cycles of individual LLMs, we assume that the LLMs remain static within the simulation, and leave the effects of their evolution for future research. For details of the prompts and post-processing steps please refer to Appendix A.8 and A.3.
@@ 4 Experiment
Datasets and Metrics. We conduct experiments on commonly used ODQA datasets, including NQ (Kwiatkowski et al., 2019), WebQ (Berant et al., 2013), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022). We preprocess the datasets following Yu et al. (2023) and Zhang et al. (2023). Given the constraints on experimental resources, we randomly select 200 samples from each test set for a comprehensive analysis. For evaluating the retrieval phase in ODQA tasks during the simulation, we utilize the widely accepted metrics of Acc@5 and Acc@20 following Karpukhin et al. (2020). These metrics assess the proportion of questions where the correct answers appear in the top 5 or top 20 retrieval results, respectively. For the answer quality of the LLM output for each iteration, we follow Chen et al. (2023) by applying the Exact Match (EM) metric, which checks if the correct answer is fully contained within the generated text. Furthermore, in Section 5, we adopt a holistic perspective to examine the RAG pipeline, with a focus on the interaction between the retrieval and generation phases and how the ranking of human-generated texts changes over time.
Retrieval and Re-ranking Methods. In our experiments, we employ a variety of retrieval methods, including the sparse method BM25, the contrastive learning-based dense retriever Contriever (Izacard et al., 2022), the advanced BGE-Base (Xiao et al., 2023) retriever, and the LLM-Embedder (Zhang et al., 2023) designed for LLMs. For the results retrieved using BM25 and BGE-Base, we separately apply the T5-based (Raffel et al., 2020) re-ranking model MonoT5-3B (Nogueira et al., 2020), the UPR-3B (Sachan et al., 2022) which uses the unsupervised capabilities of T0-3B (Sanh et al., 2022), and the BGE-Reranker (Xiao et al., 2023), which is based on the XLM-RoBERTa-Large (Conneau et al., 2020).
Generative Models. Considering the complexity and variability of real-world environments, the text that is continuously integrated into the system may be generated by a variety of LLMs. Our iterative experiments incorporate text produced by a suite of prevalent LLMs. These include GPT-3.5-Turbo (OpenAI, 2022), LLaMA2-13B-Chat (Touvron et al., 2023), Qwen-14B-Chat (Bai et al., 2023), Baichuan2-13B-Chat (Yang et al., 2023), and ChatGLM3-6B (Du et al., 2022). This enables the RAG systems to blend varied linguistic styles and knowledge, leading to results that more closely replicate real-world scenarios.
Implementation. Please refer to Appendix A.4.
@@
Figure 2: Short-Term QA performance. For each retrieval method, we document both the average performance and the range of variation exhibited by five LLMs. A red dashed line symbolizes the average EM score for zero-shot question generation by LLMs. "Ori." and "+LLMZ" represent the average EM values when models use the original dataset or a dataset enhanced with LLM-generated texts as context, respectively. Retrieval methods are abbreviated: "Contri" for Contriever, "LLM-E" for LLM-Embedder, and "BGE-B" for BGEbase.
@@ 5 Results
In this section, within the simulation framework, we meticulously examine both the initial and the extended iterations. We define the short-term effect as the immediate effects observed in the first iteration, while the long-term effect is analyzed based on the second to the tenth iteration. We investigate if the "Spiral of Silence" effect occurs within these stages and how the RAG systems might respond to it. Under the task settings of ODQA, we analyze the potential influence of the "Spiral of Silence" on RAG systems’ response patterns.
@@ 5.1 Short-Term Effects on RAG Performance
When comparing RAG system results using different retrieval methods on the original dataset versus the augmented one in the first iteration, we observe that: 1) Immediate Impact of LLM-Generated Text on the RAG System: The introduction of a minimal amount of LLM-generated text produces immediate effects on both retrieval and QA performance of the RAG system, as shown in Table 1 and Figure 2. 2) LLM-Generated Text Generally Improves Retrieval Accuracy: Table1 reveals that adding LLM-generated responses to a dataset typically enhances the accuracy of retrieval systems, as measured by Acc@5 and Acc@20 metrics. For example, using the BM25 on TriviaQA resulted in accuracy improvements of 31.2% and 19.1% respectively. However, a slight decline in Acc@5 is also observed in certain cases. This suggests a primarily positive, yet complex, impact of LLM-generated text on retrieval accuracy.
@@
3) The impact on QA performance is mixed: Due to space constraints, we only present the results of four retrieval methods. As shown in Figure 2, while the RAG system’s QA performance typically surpasses the zero-shot LLM outputs, the addition of LLM text can either enhance or impair QA performance depending on the dataset and retrieval strategy. It appears to enhance performance for TriviaQA, but for NQ and PopQA, the effect is detrimental with non-BM25 retrieval methods, suggesting that without significant retrieval enhancement, LLM text inclusion might be counterproductive.
@@ 5.2 Long-term Effects on RAG Performance
In this section, we investigate whether the shot-term effects are predictive of the long-term behavior of the system. We present the results on NQ and PopQA in Figure 3; for results on other datasets, please refer to Figure 8 in Appendix A.5. We observe that: 1) Decreased Retrieval Effectiveness Over Time: Figures 3(a) and 3(b) show a general decline in Acc@5 across successive iterations for most methods, with an average drop of 21.4% for NQ and 19.4% for PopQA from the first iteration to the last, except for a temporary improvement in BM25 during the second iteration on PopQA. This trend signals that the retrieval quality boost provided by LLM-generated text may be transient, with a propensity for degradation over time. 2) Stability in QA Performance Despite Retrieval Decline: Contrary to expectations, the QA performance does not mirror the retrieval accuracy’s decrease. As shown in Figure 3(c) and 3(d), the EM exhibit slight variations but generally maintain their level throughout the iterations. While a diminished retrieval accuracy intuitively seems to undermine the system’s capacity to output correct answers, this does not unequivocally translate into a decline in QA efficacy. In subsequent sections, we will delve deeper into the reasons behind these observations and examine the complex dynamic relationship that may exist between retrieval and QA performance.
@@ 5.3 Spiral of Silence
In the context of LLM-augmented RAG systems, we have observed a rapid shift in response to the integration of LLM-generated text, a decline in retrieval performance over time, and stability in QA performance despite retrieval decline. To explain these phenomena, we draw on the theory of the "Spiral of Silence" as posited by Noelle-Neumann (Noelle-Neumann, 1974), extending its principles to the behavior of RAG systems enhanced by LLMs. To explore the presence of a "Spiral of Silence" phenomenon, we propose three Hypotheses for investigation. (H1): Dominance of LLM-Generated Texts. Retrieval models are more likely to prioritize LLM-generated text in search results, which could result in LLM-generated text taking a dominant position in the retrieval hierarchy. (H2): Marginalization of Human-Generated Content. If human-authored text consistently loses ranking prominence through successive iterations, it may be excluded from the top results until it becomes invisible, thus creating a silence. (H3): Homogenization of Opinions. The preferential ranking of LLM-generated text could culminate in a uniformity of displayed perspectives by the RAG system, potentially sidelining the accuracy or variety of the information.
@@
Table 2: Percentage of LLM-generated documents occupying the top-5 retrieval results, after augmenting each query with five LLM-generated documents. The blue background indicates a majority presence of human-generated documents, while the purple background denotes a predominance of LLM-generated documents.
To verify (H1), we analyze Iteration1 where LLM-generated texts are first introduced to the retrieval system. We calculate the proportion of these texts appearing in the top five search results:
@@
where $c_{q}^{LLM}$ is the count of LLM-generated texts and $c_{q}^{Human}$ is the count of human-generated texts in the top five search results for query $q$. Table 2 reveals that, even with a modest inclusion of LLM-generated texts, most retrieval models often rank them at the top. This behavior supports the findings of  Dai et al. (2023a), where LLM-rewritten texts are preferred by retrieval models over the originals. Our study extends this by directly generating query-specific texts with LLMs. The preference might stem from inherent biases within the system or the actual relevance of the LLM-produced content. This suggests retrieval systems tend to favor LLM-generated texts, making them more prominent in search results, which can rapidly influence an RAG system’s behavior.
@@
Figure 5: 3-gram Self-BLEU score for the top 5 search results over iterations, from the original dataset (Ori.) to the inclusion and subsequent iterations with LLM-generated texts. For results on WebQ and TriviaQA, please refer to Figure 10 in Appendix A.5.
@@
To validate (H2), we incorporate a temporal dimension, observing the percentage change of texts generated by various LLMs and humans within the top 50 search results across different datasets over time. As shown in Figure 4, after ten iterations, the percentage of human-generated texts significantly decreased, falling below 10% for all datasets. This pattern suggests a sustained diminishing impact of human-contributed texts and hints at the possibility of their eventual exclusion from search results if trends continue.
@@
Figure 6: Correlation between the number of top 5 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by LLMs on the NQ dataset. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
@@
To explore (H3), we examine the risk of potential viewpoint homogenization in the RAG system from both the diversity and accuracy dimensions during the simulation. Diversity is quantified using the Self-BLEU score, where a higher score indicates less diversity, suggesting a convergence of viewpoints. As shown in Figure 5, upon introducing zero-shot LLM-generated texts (Iteration1), the Self-BLEU scores across different datasets experience varying degrees of change. However, over more iterations, the Self-BLEU scores for the top 5 results consistently rise and plateau across all datasets, indicating a significant reduction in textual diversity with each iterative cycle. Subsequently, we assessed whether the accuracy of the top documents returned by the IR system tends toward uniformity over time. Figure 6 charts the number of documents with the correct answer in the top 5 results ("Context Right Num") against the number of queries LLM answers correctly or incorrectly, across the Iteration1, 2, 5, 10. For simplicity, we showcase only the NQ dataset’s averaged outcomes, but we find the same trends across other datasets. It indicates that in the initial iterations, fewer correct answer documents (e.g., "Context Right Num" of 0, 1, or 2) typically correlate with a greater number of LLM-answered queries being incorrect (EM=0). Despite this, a significant fraction of the top documents still include the correct answer. When the LLM correctly answers a query (EM=1), the correct answer documents within the top results can range anywhere from 1 to 5. As the iterations continue, the frequency of having 1 to 4 correct answer documents in the top 5 results for each query diminishes, and by the Iteration10, contexts for LLM-correct queries almost always contain the correct answer, whereas contexts for LLM-incorrect answers almost always do not. This pattern demonstrates a trend towards polarization and uniformity in the accuracy of provided contexts as the RAG system iterates.
@@
At this point, we have confirmed through our experiments the presence of the "Spiral of Silence" phenomenon, as outlined in three tested hypotheses. Moreover, the dashed lines in Figure 6 represent the total number of correct and incorrect LLM answered queries, along with the Acc@5 retrieval metric over various iterations. The LLM’s rate of correct answers remains constant through the iterations, aligning to Section 5.2. However, as iterations advance, correct answers diminish within top documents for LLM-incorrect queries, reducing their contribution to the Acc@5 and thus decreasing retrieval performance. In contrast, for LLM-correct queries, more retrieved documents containing the correct answer do not affect Acc@5 or EM. This pattern is explicable by the "Spiral of Silence" theory discussed in Section 5.2, which accounts for the observed dip in IR results and the sustained QA performance.
5.4 Effects of Spiral of Silence on ODQA

### v3

@@ Abstract
The practice of Retrieval-Augmented Generation (RAG), which integrates Large Language Models (LLMs) with retrieval systems, has become increasingly prevalent. However, the repercussions of LLM-derived content infiltrating the web and influencing the retrieval-generation feedback loop are largely uncharted territories. In this study, we construct and iteratively run a simulation pipeline to deeply investigate the short-term and long-term effects of LLM text on RAG systems. Taking the trending Open Domain Question Answering (ODQA) task as a point of entry, our findings reveal a potential digital “Spiral of Silence” effect, with LLM-generated text consistently outperforming human-authored content in search rankings, thereby diminishing the presence and impact of human contributions online. This trend risks creating an imbalanced information ecosystem, where the unchecked proliferation of erroneous LLM-generated content may result in the marginalization of accurate information. We urge the academic community to take heed of this potential issue, ensuring a diverse and authentic digital information landscape. [^1] We release the resources at https://github.com/VerdureChen/SOS-Retrieval-Loop
@@ 1 Introduction
Figure 1: The evolution of RAG systems after introducing LLM-generated texts, where the “Spiral of Silence” effect gradually emerges.
@@
Further examination reveals a bias in search systems towards LLM-generated texts, which consistently rank higher than human-written content. As LLM-generated texts increasingly dominate the search results, the visibility and influence of human-authored web content diminish, fostering a digital “Spiral of Silence” effect. This effect aptly explains what we observe in our simulations and reveals the potential negative impact of LLM-generated texts on the information ecosystem: while LLM-generated texts sometimes provide a more effective IR experience in the short term, in the long term they may lead to the invisibility of human-authored content, the homogenization of search results, and the inaccessibility of certain accurate information, thereby adversely affecting public knowledge acquisition and decision-making.
The contributions of this paper are threefold: 1) We propose an iterative pipeline to investigate the short-term and long-term impacts of LLM-generated text on RAG systems. 2) We study the potential emergence of a “Spiral of Silence” phenomenon within RAG systems. 3) We analyze the implications of this phenomenon, offering a new perspective on the dynamic interplay between LLM-generated content and RAG systems.
@@ 2 Related Works
Spiral of Silence. The “Spiral of Silence” theory (Noelle-Neumann, 1974), is a seminal theory within the field of communications that describes how people may suppress their views to avoid isolation, thus often reinforcing dominant public opinions (Scheufle and Moy, 2000; Liu et al., 2019; Lin et al., 2022). We shift focus to a novel “passive human silence” influenced by LLMs, where rapid AI content production and biased search algorithms potentially marginalize human contributions in public discourse. This theory stands apart from concepts such as “echo chambers", “filter bubbles", and “degenerate feedback loops” prevalent in recommendation systems  (Alatawi et al., 2021; Chitra and Musco, 2020; Jiang et al., 2019). While these terms describe the narrowing of informational scope as users engage with algorithmic systems, the “Spiral of Silence” theory proposes a scenario where human users are compelled into silence in public discourse due to the influence of LLMs and IR systems. This phenomenon goes beyond mere selective exposure or algorithmic recommendations. Our study explores how LLM-generated text might induce a “Spiral of Silence” in RAG systems over time. For further discussion regarding the rationale for applying this theory to our study, please see Appendix A.1.
@@ 3.1 Preliminaries
Generation Stage, executed by the generation function $G:(P\times Q\times D^{\prime}\times K)\rightarrow S$. Its task is to utilize the prompt $p\in P$, the query $q\in Q$, the related document subset $D^{\prime}$, and the knowledge of LLMs $K$ to construct the answer $s$.
@@ 3.2 Simulation Process
Our simulation process starts with a pure human-authored text dataset and gradually introduces the LLM-generated text, observing how this change over time affects the RAG system. Adhering to the specifications outlined in Section 3.1, the RAG architecture is instantiated and expanded with additional details. In the retrieval stage, we apply sparse and dense retrieval strategies to obtain a candidate document set that is relevant to the query. Additionally, we also have the option to perform a re-ranking of the candidate documents to further optimize the process. In the generation stage, we use the LLMs which are widely used to generate responses. To accurately simulate the evolution process of the RAG system, we specifically use an iteratively updated indexing structure that supports incorporating the newly generated LLM text into the index in each iteration, keeping the dataset updated for subsequent retrieval and evaluation.
@@
The pseudo-code for this process is presented in Appendix A.2. Through the simulation process, we observe how LLM-generated text influences the RAG systems and how this impact evolves with data accumulation. While the main simulation assumes that the LLMs remain static due to their relatively infrequent update cycles, we also conduct experiments on the effects of LLM evolution over time in Appendix A.6. For prompt and post-processing details, see Appendix A.9 and A.3.
@@ 4 Experiment
Datasets and Metrics. We conduct experiments on commonly used ODQA datasets, including NQ (Kwiatkowski et al., 2019), WebQ (Berant et al., 2013), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022). We preprocess the datasets following Yu et al. (2023) and Zhang et al. (2023). Given the constraints on experimental resources, we randomly select 200 samples from each test set. When evaluating the retrieval phase, we utilize Acc@5 and Acc@20 following Karpukhin et al. (2020). These metrics assess the proportion of questions where the correct answers appear in the top 5 or top 20 retrieval results, respectively. For the answer quality of the LLM output for each iteration, we follow Chen et al. (2023) by applying the Exact Match (EM) metric, which checks if the correct answer is fully contained within the generated text. Furthermore, in Section 5, we adopt a holistic perspective to examine the RAG pipeline, with a focus on the interaction between the retrieval and generation phases and how the ranking of human-generated texts changes over time.
Retrieval and Re-ranking Methods. In our experiments, we employ a variety of retrieval methods, including the sparse model BM25, the contrastive learning-based dense retriever Contriever (Izacard et al., 2022), the advanced BGE-Base (Xiao et al., 2023) retriever, and the LLM-Embedder (Zhang et al., 2023) designed for LLMs. For the results retrieved using BM25 and BGE-Base, we separately apply the T5-based (Raffel et al., 2020) re-ranking model MonoT5-3B (Nogueira et al., 2020), the UPR-3B (Sachan et al., 2022) which uses the unsupervised capabilities of T0-3B (Sanh et al., 2022), and the BGE-Reranker (Xiao et al., 2023), which is based on the XLM-RoBERTa-Large (Conneau et al., 2020).
Generative Models. Considering the complexity and variability of real-world environments, the text that is continuously integrated into the system may be generated by a variety of LLMs. Our iterative experiments incorporate text produced by a suite of prevalent LLMs. These include GPT-3.5-Turbo (OpenAI, 2022), LLaMA2-13B-Chat (Touvron et al., 2023), Qwen-14B-Chat (Bai et al., 2023), Baichuan2-13B-Chat (Yang et al., 2023), and ChatGLM3-6B (Du et al., 2022). This enables the RAG systems to blend varied linguistic styles and knowledge, leading to results that more closely replicate real-world scenarios. For more implementation details, please refer to Appendix A.4.
@@
Figure 2: Short-Term QA performance. For each retrieval method, we present both the average performance and the range of variation exhibited by five LLMs. A red dashed line symbolizes the average EM score for zero-shot question generation by LLMs. “Ori.” and “+LLMZ” represent the average EM values when models use the original dataset or a dataset enhanced with LLM-generated texts as context, respectively. Retrieval methods are abbreviated: “Contri” for Contriever, “LLM-E” for LLM-Embedder, and “BGE-B” for BGEbase.
@@ 5 Results
In this section, we examine both initial and extended iterations within the simulation framework. We define the short-term effect as the immediate effects observed in the first iteration, while the long-term effect is analyzed from the second to the tenth iteration. We investigate the occurrence of the “Spiral of Silence” effect and how RAG systems respond. Under the task settings of ODQA, we analyze the potential influence of the “Spiral of Silence” on RAG systems’ response patterns.
@@ 5.1 Short-Term Effects on RAG Performance
When comparing RAG system results using different retrieval methods on the original dataset versus the augmented one in the first iteration, we observe that: 1) Immediate Impact of LLM-Generated Text on the RAG System: The introduction of a minimal amount of LLM-generated text produces immediate effects on both retrieval and QA performance of the RAG system, as shown in Table 1 and Figure 2. Specifically, both retrieval and generation performance exhibit noticeable fluctuations. These changes highlight the sensitivity of the system to even small modifications made by LLM-generated text. 2) LLM-Generated Text Generally Improves Retrieval Accuracy: Table 1 reveals that adding LLM-generated responses to a dataset typically enhances the accuracy of retrieval systems, as measured by Acc@5 and Acc@20 metrics. For example, using the BM25 on TriviaQA resulted in accuracy improvements of 31.2% and 19.1% respectively. However, a slight decline in Acc@5 is also observed in certain cases. This suggests a primarily positive, yet complex, impact of LLM-generated text on retrieval accuracy. 3) The Impact on QA Performance is Mixed: Due to space constraints, we only present the results of four retrieval methods. As shown in Figure 2, while the RAG system’s QA performance typically surpasses the zero-shot LLM outputs, the addition of LLM text can either enhance or impair QA performance depending on the dataset and retrieval strategy. It appears to enhance performance for TriviaQA, but for NQ and PopQA, the effect is detrimental with non-BM25 retrieval methods, suggesting that without significant retrieval enhancement, LLM text inclusion might be counterproductive.
@@ 5.2 Long-term Effects on RAG Performance
In this section, we investigate whether the short-term effects are predictive of the long-term behavior of the system. We present the results on NQ and PopQA in Figure 3. For results on other datasets, please refer to Figure 8 in Appendix A.5, where we observe consistent patterns across these datasets. We find that: 1) Decreased Retrieval Effectiveness Over Time: Figures 3(a) and 3(b) show a general decline in Acc@5 across successive iterations for most methods, with an average drop of 21.4% for NQ and 19.4% for PopQA from the first iteration to the last, except for a temporary improvement in BM25 during the second iteration on PopQA. This trend signals that the retrieval quality boost provided by LLM-generated text may be transient, with a propensity for degradation over time. 2) Stability in QA Performance Despite Retrieval Decline: Contrary to expectations, the QA performance does not mirror the retrieval accuracy’s decrease. As shown in Figure 3(c) and 3(d), the EM exhibit slight variations but generally maintain their level throughout the iterations. While a diminished retrieval accuracy intuitively seems to undermine the system’s capacity to output correct answers, this does not unequivocally translate into a decline in QA efficacy. In subsequent sections, we will delve deeper into the reasons behind these observations and examine the complex dynamic relationship that may exist between retrieval and QA performance.
@@ 5.3 Spiral of Silence
In the context of LLM-augmented RAG systems, we have observed a rapid shift in response to the integration of LLM-generated text, a decline in retrieval performance over time, and stability in QA performance despite retrieval decline. To explain these phenomena, we draw on the theory of the “Spiral of Silence” as posited by Noelle-Neumann (1974), extending its principles to the behavior of RAG systems enhanced by LLMs. To explore the presence of a “Spiral of Silence” phenomenon, we propose three Hypotheses for investigation. (H1): Dominance of LLM-Generated Texts: Retrieval models are more likely to prioritize LLM-generated text in search results, which could result in LLM-generated text taking a dominant position in the retrieval hierarchy. (H2): Marginalization of Human-Generated Content: If human-authored text consistently loses ranking prominence through successive iterations, it may be excluded from the top results until it becomes invisible, thus creating silence. (H3): Homogenization of Opinions: The preferential ranking of LLM-generated text could culminate in a uniformity of displayed perspectives by the RAG system, potentially sidelining the accuracy or variety of the information.
@@
Table 2: Percentage of LLM-generated documents occupying the top 5 retrieval results, after augmenting each query with five LLM-generated documents. The blue background indicates a majority presence of human-generated documents, while the purple background denotes a predominance of LLM-generated documents.
To verify (H1), we analyze Iteration 1 where LLM-generated texts are first introduced to the retrieval system. We calculate the proportion of these texts appearing in the top 5 search results:
@@
where $c_{q}^{LLM}$ is the count of LLM-generated texts and $c_{q}^{Human}$ is the count of human-generated texts in the top 5 search results for query $q$. Table 2 reveals that, even with a modest inclusion of LLM-generated texts, most retrieval models often rank them at the top. This behavior supports the findings of  Dai et al. (2023a), where LLM-rewritten texts are preferred by retrieval models over the originals. Our study extends this by directly generating query-specific texts with LLMs. The preference might stem from inherent biases within the system or the actual relevance of the LLM-produced content. This suggests retrieval systems tend to favor LLM-generated texts, making them more prominent in search results, which can rapidly influence an RAG system’s behavior.
@@
Figure 5: 3-gram Self-BLEU score for the top 5 search results over iterations, from the original dataset (Ori.) to subsequent iterations including LLM-generated texts. For results on WebQ and TriviaQA, please refer to Figure 10 in Appendix A.5.
@@
To validate (H2), we incorporate a temporal dimension, observing the percentage change of texts generated by various LLMs and humans within the top 50 search results across different datasets over time. As shown in Figure 4, after ten iterations, the percentage of human-generated texts significantly decreased, falling below 10% for all datasets. This pattern suggests a sustained diminishing impact of human-contributed texts and hints at the possibility of their eventual exclusion from search results if the trend continues.
@@
Figure 6: Correlation between the number of top 5 search results containing the correct answer (“Context Right Num”) and the accuracy of responses given by LLMs on the NQ dataset. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval, and ranking methods.
@@
To explore (H3), we examine the risk of potential viewpoint homogenization in the RAG system from both the diversity and accuracy dimensions during the simulation. Diversity is quantified using Self-BLEU  (Zhu et al., 2018), which works by comparing a generated text with other generated texts to measure similarity. High similarity results in a high Self-BLEU score, indicating lower diversity and suggesting a convergence of viewpoints. As shown in Figure 5, upon introducing zero-shot LLM-generated texts (Iteration 1), the Self-BLEU scores across different datasets experience varying degrees of change. However, over more iterations, the Self-BLEU scores for the top 5 results consistently rise and plateau across all datasets, indicating a significant reduction in textual diversity with each iterative cycle. Subsequently, we assessed whether the accuracy of the top documents returned by the IR system tends toward uniformity over time. Figure 6 charts the number of documents with the correct answer in the top 5 results ("Context Right Num") against the number of queries LLM answers correctly or incorrectly, across the Iteration1, 2, 5, 10. For simplicity, we showcase only the NQ dataset’s averaged outcomes, but we find the same trends across other datasets. It indicates that in the initial iterations, fewer correct answer documents (e.g., “Context Right Num” of 0, 1, or 2) typically correlate with a greater number of LLM-answered queries being incorrect (EM=0). Despite this, a significant fraction of the top documents still include the correct answer. When the LLM correctly answers a query (EM=1), the correct answer documents within the top results can range anywhere from 1 to 5. As the iterations continue, the frequency of having 1 to 4 correct answer documents in the top 5 results for each query diminishes, and by the Iteration10, contexts for LLM-correct queries almost always contain the correct answer, whereas contexts for LLM-incorrect answers almost always do not. This pattern demonstrates a trend towards polarization and uniformity in the accuracy of provided contexts as the RAG system iterates.
@@
At this point, we have confirmed through our experiments the presence of the “Spiral of Silence” phenomenon, as outlined in three tested hypotheses. Moreover, the dashed lines in Figure 6 represent the total number of correct and incorrect LLM answered queries, along with the Acc@5 retrieval metric over various iterations. The LLM’s rate of correct answers remains constant through the iterations, aligning to Section 5.2. However, as iterations advance, correct answers diminish within top documents for LLM-incorrect queries, reducing their contribution to the Acc@5 and thus decreasing retrieval performance. In contrast, for LLM-correct queries, more retrieved documents containing the correct answer do not affect Acc@5 or EM. Thus, the pattern discussed in Section 5.2 can be explained by the “Spiral of Silence” theory, which accounts for the observed dip in IR results and the sustained QA performance.
5.4 Effects of “Spiral of Silence” on ODQA